### v2

@@ 4.3 Deriving a Final Feature Set
Figure 8: Comparison of feature set performance across dataset versus dataset (DD) cases. The left side of the heatmap displays the results of applying feature sets obtained from the first step of the Genetic Algorithm (GA) to all DD data, yielding 64 results. On the right, the performance of the features grouped according to their frequency, focusing on the intersection of features obtained from the output of the GA algorithm.
@@ 5.1 Model Selection
Figure 9: F1 scores and average inference times of various ML algorithms applied to DD datasets. RF and XGB demonstrate the highest F1 scores of 0.799 and 0.780, respectively, while DT show the fastest inference time.
@@ 5.2 Evaluating Model Generalizability
To test our hypothesis about the fragility of statistical features, we needed to compare our packet-based approach against strong, representative baselines for flow-based and window-based methods. For this purpose, we selected CICFlowmeter [38] and Kitsune [39] not as published methods to compare against, but as well-established tools for extracting comprehensive sets of flow and window-based features, respectively. These tools are powerful and generate a wide array of statistical features, making them ideal for creating robust baseline models that effectively represent the statistical methods prevalent in the DI literature. We then applied our own rigorous feature selection and evaluation pipeline to these feature sets to ensure a fair and direct comparison of the feature types themselves.
To ensure a fair comparison, feature and model selection were repeated independently (see Appendix Sections B and C for results). We also compare against IoTDevID [13], another packet header-based DI method that aimed to build generalizable DI models and demonstrated improved performance over alternative methods.
@@
Table 2: F1 scores per device in all 8 DD cases for MonIoTr data, with green for higher and red for lower success.
@@ 5.3 Feature Usage
Notably, IoTDevID’s feature set is rich in features related to BOOTP, DNS, and EAPOL protocols, whereas GeMID lacks these. This may be due to a dataset-specific bias, owing to the reliance on a single dataset—Aalto—for IoTDevID’s feature selection process. This is supported by the observation that while GeMID outperforms IoTDevID in experiments using the UNSW and MonIoTr datasets, IoTDevID shows markedly better performance when applied to the Aalto dataset [13] (see Appendix Table 8).
@@
Another reason for the difference in performance may be the larger initial pool of extracted features in GeMID (332 features) compared to IoTDevID (112 features). Although both studies generally use the same protocols, GeMID captures more sub-features within each protocol, providing finer-grained information in the feature set. In terms of selected features, GeMID offers more comprehensive protocol coverage, starting with a richer set of HTTP features and extending to broader coverage of TCP and UDP protocols. This includes features such as sequence numbers, timing metrics, and window size information, which enhance its ability to capture more detailed network behavior (see Figure 7). Unlike in the IoTDevID study, we retained features containing session-based identifiers (such as TCP acknowledgment, TCP sequence, and source/destination port numbers). Whilst these can lead to over-optimistic metric scores [40], this is only the case when sessions are split between test and train sets, e.g., when using CV. When this is controlled for using stricter evaluation contexts (as in this study), the inclusion of these features can enhance model performance by capturing meaningful relationships between packets.
@@ 7 Conclusions
Comparative analysis of DI methods showed that packet header-based features performed better than network statistics and window methods for building generalizable DI models. Packet features offer a more consistent and reliable basis for model training because they are less influenced by network-specific variables. This claim is most strongly supported by our generalization gap analysis (Table 1), where the performance of statistical methods like CICFlowMeter and Kitsune degraded by over 100% more than our packet-based GeMID method when moving from single-environment validation to cross-environment testing (CV-DD $\Delta$F1 of 0.393 and 0.464 vs. 0.185). This demonstrates that packet features offer a more consistent and reliable basis for model training because they capture device-intrinsic attributes rather than network-specific variables. This finding highlights the importance of selecting features that inherently capture device-specific attributes.
@@
Our findings underscore that statistical methods, despite their prevalence, are unreliable for DI due to their entanglement with network conditions. This challenges nearly half of existing research [18, 19], urging a re-evaluation of studies reliant on flow or window statistics. Future work should prioritize packet-based or similarly device-centric features to ensure generalizability, reshaping IoT security practices.
@@ Appendix B CICFlowmeter Feature Selection Process and Results
Figure 10: Comparison of kappa scores for CICFlowmeter features in UNSW datasets using CV (blue) and isolated methods, SS (red) and DD (green). The CV method tends to overestimate feature utility, showing higher scores for many attributes, while SS and DD produce more realistic evaluations.
Figure 11: A voting system for CICFlowmeter features that gives one vote to each feature with a kappa value different from 0 that is individually predictive within CV (blue), SS (red) & DD (green).
@@
Figure 13: Comparison of CICFlowmeter feature set performance across dataset versus dataset (DD) cases. The left side of the heatmap displays the results of applying feature sets obtained from the first step of the Genetic Algorithm (GA) to all DD data, yielding 64 results. On the right, the performance of the features grouped according to their frequency, focusing on the intersection of features obtained from the output of the GA algorithm.
@@ Appendix C Kitsune Feature Selection Process and Results
Figure 15: Comparison of kappa scores for Kitsune features in UNSW datasets using CV (blue) and isolated methods, SS (red) and DD (green). The CV method tends to overestimate feature utility, showing higher scores for many attributes, while SS and DD produce more realistic evaluations. This discrepancy highlights the potential for information leakage in cross-validation and the importance of using isolated validation methods for assessing feature utility in ML-based DI models.
Figure 16: A voting system for Kitsune features that gives one vote to each feature with a kappa value different from 0 that is individually predictive within CV (blue), SS (red) & DD (green).
@@
Figure 18: Comparison of Kitsune feature set performance across dataset versus dataset (DD) cases. The left side of the heatmap displays the results of applying feature sets obtained from the first step of the Genetic Algorithm (GA) to all DD data, yielding 64 results. On the right, the performance of the features grouped according to their frequency, focusing on the intersection of features obtained from the output of the GA algorithm.
@@ E.3 Results from Aggregation Algorithm
Table 10: GeMDID F1 scores per device in all 8 DD cases for MonIoTr data with aggregation algorithm.
@@ E.4 Hyperparameter Optimization
NB Var Smoothing $10^{0}$ $10^{-9}$ $1.52\times 10^{-6}$ $8.11\times 10^{-9}$ $2.31\times 10^{-6}$ $8.11\times 10^{-9}$
LR C $10^{-5}$ 100

### v3

@@ 4.3 Deriving a Final Feature Set
Figure 8: Comparison of feature set performance across dataset versus dataset (DD) cases. The left side of the heatmap displays the results of applying feature sets obtained from the first step of the Genetic Algorithm (GA) to all $D$ data, yielding 64 results. On the right, the performance of the features grouped according to their frequency, focusing on the intersection of features obtained from the output of the GA algorithm.
@@ 5.1 Model Selection
Figure 9: F1 scores and average inference times of various ML algorithms applied to $D$ datasets. RF and XGB demonstrate the highest F1 scores of 0.799 and 0.780, respectively, while DT show the fastest inference time.
@@ 5.2 Evaluating Model Generalizability
To test our hypothesis about the fragility of statistical features, we needed to compare our packet-based approach against strong, representative baselines for flow-based and window-based methods. Rather than attempting to replicate individual studies, which often rely on different evaluation criteria and settings, we chose a more fundamental approach. For this purpose, we selected CICFlowmeter [flowmeter] and Kitsune [mirsky2018kitsune] not as published methods to compare against, but as well-established tools for extracting broad sets of flow- and window-based features, respectively (see Table 1). These tools generate feature sets that cover the majority of statistical features commonly used across the DI literature, making them suitable for constructing baseline models that fairly represent these entire classes of methods. We then applied our own rigorous feature selection and evaluation pipeline to these feature sets to ensure a controlled, fair, and direct comparison of the feature types themselves. We also compare against IoTDevID [kostas22IoTDevID], another packet header-based DI method that aimed to build generalizable DI models and demonstrated improved performance over alternative methods (see Appendix Sections B and C for results).
@@
Table 2: F1 scores per device in all 8 $D$ cases for MonIoTr data, with green for higher and red for lower success.
@@ 5.3 Feature Usage
Notably, IoTDevID’s feature set is rich in features related to BOOTP, DNS, and EAPOL protocols, whereas GeMID lacks these. This may be due to a dataset-specific bias, owing to the reliance on a single dataset—Aalto—for IoTDevID’s feature selection process. This is supported by the observation that while GeMID outperforms IoTDevID in experiments using the UNSW and MonIoTr datasets, IoTDevID shows markedly better performance when applied to the Aalto dataset [kostas22IoTDevID] (see Appendix Table 8).
@@
Another reason for the difference in performance may be the larger initial pool of extracted features in GeMID (332 features) compared to IoTDevID (112 features). Although both studies generally use the same protocols, GeMID captures more sub-features within each protocol, providing finer-grained information in the feature set. In terms of selected features, GeMID offers more comprehensive protocol coverage, starting with a richer set of HTTP features and extending to broader coverage of TCP and UDP protocols. This includes features such as sequence numbers, timing metrics, and window size information, which enhance its ability to capture more detailed network behavior (see Figure 7). Unlike in the IoTDevID study, we retained features containing session-based identifiers (such as TCP acknowledgment, TCP sequence, and source/destination port numbers). Whilst these can lead to over-optimistic metric scores [kostas2024individual], this is only the case when sessions are split between test and train sets, e.g., when using CV. When this is controlled for using stricter evaluation contexts (as in this study), the inclusion of these features can enhance model performance by capturing meaningful relationships between packets.
@@ 6.2 Study Limitations and Future Work
A further consideration is the generalizability of the feature selection process itself. Our methodology deliberately used the UNSW dataset family for feature and model selection and the entirely independent MonIoTr dataset family for final validation. This design ensures that our final performance metrics are not biased by the selection process. While we believe that our process, and the datasets used, provide strong evidence for the robustness of our framework, we encourage others to investigate enhancements to our model, such as by studying other datasets for feature selection and testing. Though our UNSW feature-based selection and testing on MonIoTr demonstrate that our model selected fundamental, device-intrinsic properties (and are not dataset-specific), it is likely that the exact features used would vary if selected on a different device population, due to both the different network behaviors of the devices and the stochastic nature of the genetic algorithm used. Nevertheless, our framework is designed to be a robust process for identifying the most stable features in any given context, and we anticipate that a core set of highly predictive features would remain consistent across datasets.
@@ 7 Conclusions
[…]
@@
Our findings underscore that statistical methods, despite their prevalence, are unreliable for DI due to their entanglement with network conditions. This challenges nearly half of existing research [kostas23thesis, jmila2022survey], urging a re-evaluation of studies reliant on flow or window statistics. Future work should prioritize packet-based or similarly device-centric features to ensure generalizability, reshaping IoT security practices.
@@ Appendix B CICFlowmeter Feature Selection Process and Results
Figure 10: Comparison of kappa scores for CICFlowmeter features in UNSW datasets using CV (blue) and isolated methods, SS (red) and $D$ (green). The CV method tends to overestimate feature utility, showing higher scores for many attributes, while SS and $D$ produce more realistic evaluations.
Figure 11: A voting system for CICFlowmeter features that gives one vote to each feature with a kappa value different from 0 that is individually predictive within CV (blue), SS (red) & $D$ (green).
@@
Figure 13: Comparison of CICFlowmeter feature set performance across dataset versus dataset (DD) cases. The left side of the heatmap displays the results of applying feature sets obtained from the first step of the Genetic Algorithm (GA) to all $D$ data, yielding 64 results. On the right, the performance of the features grouped according to their frequency, focusing on the intersection of features obtained from the output of the GA algorithm.
@@ Appendix C Kitsune Feature Selection Process and Results
Figure 15: Comparison of kappa scores for Kitsune features in UNSW datasets using CV (blue) and isolated methods, SS (red) and $D$ (green). The CV method tends to overestimate feature utility, showing higher scores for many attributes, while SS and $D$ produce more realistic evaluations. This discrepancy highlights the potential for information leakage in cross-validation and the importance of using isolated validation methods for assessing feature utility in ML-based DI models.
Figure 16: A voting system for Kitsune features that gives one vote to each feature with a kappa value different from 0 that is individually predictive within CV (blue), SS (red) & $D$ (green).
@@
Figure 18: Comparison of Kitsune feature set performance across dataset versus dataset (DD) cases. The left side of the heatmap displays the results of applying feature sets obtained from the first step of the Genetic Algorithm (GA) to all $D$ data, yielding 64 results. On the right, the performance of the features grouped according to their frequency, focusing on the intersection of features obtained from the output of the GA algorithm.
@@ E.3 Results from Aggregation Algorithm
Table 10: GeMDID F1 scores per device in all 8 $D$ cases for MonIoTr data with aggregation algorithm.
@@ E.4 Hyperparameter Optimization
NB Var Smoothing $10^{0}$ - $10^{-9}$ $1.52\times 10^{-6}$ $8.11\times 10^{-9}$ $2.31\times 10^{-6}$ $8.11\times 10^{-9}$
LR C $10^{-5}$ - 100